# Contracts with Inspections

### v1

@@ 2 Model
In order to introduce the agent’s expected utility under inspection scheme $(i,\alpha,p)$, we observe that if the agent takes the incentivized action (action $i$), then the agent’s utility is $\alpha f(i)-c(i)$ in expectation. If, instead, the agent decides to take action $j\neq i$, as long as the principal does not inspect a set containing either $i$ or $j$, then we say that the agent is not caught and is paid the promised $\alpha$-fraction of the reward (in case of success). Otherwise (the principal inspects some set containing $i$ or $j$), the agent is caught and thus paid 00. In expectation, we have that, for $j\neq i$, the utility is $\alpha f(j)\cdot(1-\sum_{S:\{i,j\}\cap S\neq\emptyset}p(S))-c(j)$. [^1] Notice that $1-\sum_{S:\{i,j\}\cap S\neq\emptyset}p(S)$ is precisely the probability the agent does not get caught by the principal. To summarize, given an inspection scheme $(i,\alpha,p)$ the principal proposes, the agent’s expected utility from selecting action $j$ is:
@@
The goal of the principal is to find an incentive-compatible (IC) inspection scheme $(i,\alpha,p)$, where, by incentive-compatible, we mean that $i\in\arg\max_{\ell\in A}u_{\text{a}}(i,\alpha,p,\ell)$, i.e., the agent is incentivized to take the suggested action $i$. Among the incentive-compatible inspection schemes, the principal wants to find the one with the highest utility, i.e., $u_{\text{p}}(i,\alpha,p,i)$ is maximized, and $i\in\arg\max_{\ell\in A}u_{\text{a}}(i,\alpha,p,\ell)$.
@@
[…]
@@
Note that the modification to the inspection scheme of 2.1 above preserves determinism, i.e., if the initial inspection scheme $(i,\alpha,p)$ is deterministic, then also $(i,\alpha,p^{\prime})$ created by 2.1 is.
@@
[…]
@@
Both value and demand oracles are considered standard in combinatorial optimization problems over set functions. In markets for goods (e.g., combinatorial auctions), a demand query corresponds to the best bundle to purchase given item prices. Demand oracles have proven useful in previous studies on combinatorial contracts [16] and multi-agent contracts [18].
@@ 3 Deterministic Inspection Schemes
In this section, we present a polynomial time algorithm that finds the best contract with a deterministic inspection scheme. Formally we prove the following theorem:

### v2

@@ 2 Fully-Observable Inspection Model
In order to introduce the agent’s expected utility under inspection scheme $(i,\alpha,p)$, we observe that if the agent takes the incentivized action (action $i$), then the agent’s utility is $\alpha f(i)-c(i)$ in expectation. If, instead, the agent decides to take action $j\neq i$, as long as the principal does not inspect a set containing either $i$ or $j$, then we say that the agent is not caught and is paid the promised $\alpha$-fraction of the reward (in case of success). Otherwise (the principal inspects some set containing $i$ or $j$), the agent is caught and thus paid 0. In expectation, we have that, for $j\neq i$, the utility is $\alpha f(j)\cdot(1-\sum_{S:\{i,j\}\cap S\neq\emptyset}p(S))-c(j)$. [^3] Notice that $1-\sum_{S:\{i,j\}\cap S\neq\emptyset}p(S)$ is precisely the probability the agent does not get caught by the principal. To summarize, given an inspection scheme $(i,\alpha,p)$ the principal proposes, the agent’s expected utility from selecting action $j$ is:
@@
The goal of the principal is to find an incentive-compatible (IC) inspection scheme $(i,\alpha,p)$, where, by incentive-compatible, we mean that $i\in\arg\max_{\ell\in A}u_{\text{a}}(i,\alpha,p,\ell)$, i.e., the agent is incentivized to take the suggested action $i$ (we assume that in case of a tie, the agent breaks the tie in favor of the suggested action). Among the incentive-compatible inspection schemes, the principal wants to find the one with the highest utility, i.e., $u_{\text{p}}(i,\alpha,p,i)$ is maximized, and $i\in\arg\max_{\ell\in A}u_{\text{a}}(i,\alpha,p,\ell)$.
@@
  • $j\in\arg\max_{\ell\in A}u_{\text{a}}(i,\alpha,p,\ell)\Longleftrightarrow j\in\arg\max_{\ell\in A}u_{\text{a}}(i,\alpha,p^{\prime},\ell)$;
  • $u_{\text{p}}(i,\alpha,p^{\prime},j)\geq u_{\text{p}}(i,\alpha,p,j)$.
@@
Note that the modification to the inspection scheme of Observation 2.1 above preserves determinism, i.e., if the initial inspection scheme $(i,\alpha,p)$ is deterministic, then also $(i,\alpha,p^{\prime})$ created by Observation 2.1 is.
@@
It is well known that $\text{additive}\subset\text{submodular}\subset\text{XOS}\subset\text{subadditive}$ and all containment relations are strict [35].
@@
Both value and demand oracles are considered standard in combinatorial optimization problems over set functions. In markets for goods (e.g., combinatorial auctions), a demand query corresponds to the best bundle to purchase given item prices. Demand oracles have proven useful in previous studies on combinatorial contracts [16] and multi-agent contracts [18]. We assume that all values have a description of size $O(1)$, as our goal is to understand the (computational and information) complexity of the problem as the number of actions grows large.
@@ 3 Deterministic Inspection Schemes
In this section, we present a polynomial time algorithm that finds the best contract with a deterministic inspection scheme. Formally, we prove the following theorem (see proof in Appendix A):
@@
4 Randomized Inspection Schemes: Submodular Inspection Cost
In this section, we shift our attention to randomized inspection schemes, recalling that the gap between deterministic and randomized inspection schemes in terms of the principal’s expected utility can be as large as $\Omega(n)$ (see Appendix C for further details). Our main result is a polynomial time algorithm that finds the optimal randomized inspection scheme for submodular inspection cost functions. Formally, we show the following:
Theorem 4.1.
Before describing the algorithm and showing why Theorem 4.1 holds, we present several definitions and properties of submodular functions.
Definition 4.2.
Given a function $v:2^{A}\rightarrow\mathbb{R}$, the Lovász extension of $v$ is $v^{L}:[0,1]^{A}\rightarrow\mathbb{R}$ where
$v^{L}(x)=\mathbb{E}_{\lambda\sim U[0,1]}[v(\{i\in A\mid x_{i}\geq\lambda\})].$
It is known that the Lovász extension of $v$ is the convex closure of $v$ which implies the following lemma (for which we add a formal proof in Appendix A for completeness). The next lemma states that, given the marginals of the inspection distribution $p(j)$, if the inspection cost function is submodular, then one can find the distribution with the minimum expected inspection cost consistent with those marginals.
Lemma 4.3.
  1. 1.
    $\sum_{S\subseteq A}p^{\prime}(S)=\sum_{S\subseteq A}p(S)$;
  2. 2.
  3. 3.
4.1 Algorithm
Our algorithm reformulates the problem of finding the best inspection scheme as solving a polynomial number of optimization problems, each with a constant number of variables and constraints. We first write it as a program over an exponential number of variables (the probabilities to inspect each set $p(S)$, and the payment $\alpha$). We then partition the interval $[0,1]$ of options for $\alpha$ to sub-intervals such that within each sub-interval the marginal probabilities that the actions are needed to be inspected have the same relative order. This allows us to solve the program in each sub-interval separately, and use Lemma 4.3 to decrease the number of variables from the objective function from an exponential number to a constant number. We then solve each optimization (non-linear) program, and return the best one. We now explain our algorithm formally:
Step 1: Going over the actions.
We fix action $i\in A$. If $c(i)=0$, then the inspection scheme $(i,0,p(\emptyset))$, where $p_{\emptyset}$ is the deterministic inspection distribution that inspects $\emptyset$ with probability $1$, is the optimal IC inspection scheme that incentivizes action $i$ (since the principal pays 0).
Otherwise, we have that $f(i)>c(i)>0$, then by using Observation 2.1 we can formulate the problem of finding the optimal (randomized) IC inspection scheme that incentivizes action $i$ as the following quadratic program:
$\min_{\alpha,p}\$$\alpha f(i)+\sum_{S\subseteq A}p(S)\cdot v(S)$
$\alpha\in[0,1],p\in\Delta(2^{A}),p(i)=p(\{i\}).$
We note that the constraints written above are the incentive-compatibility constraints, written in terms of marginal probabilities of inspection $p(j)=\sum_{S:j\in S}p(S)$, that we need to satisfy for the agent to take action $i$. Since $\bot\in A$, we can add a constraint to Program (2) of the form $\alpha\geq\frac{c(i)}{f(i)}$ without changing feasibility. Next, for each $j\in A$, if $f(j)=0$, the incentive-compatibility constraint is less restrictive than $\alpha\geq\frac{c(i)}{f(i)}$ and can thus be discarded. Else ($f(j)>0$), the incentive-compatibility constraint is equivalent to
$p(j)\geq 1-p(i)-\frac{\alpha f(i)-c(i)+c(j)}{\alpha f(j)}.$
Step 2: Partitioning the interval $[0,1]$ to sub-intervals.
Let us now define the set of $\alpha$’s for a specific action $i$ and each of the $n-1\choose 2$ pairs $j,j^{\prime}\neq i$ as the $\alpha$’s respectively satisfying $\eta_{j}(\alpha,p(i))=\eta_{j^{\prime}}(\alpha,p(i))$. Formally, let $C_{i}$ be the set of all $\alpha_{i}(j,j^{\prime})$’s such that $f(j)\neq f(j^{\prime})$ and that satisfy
Step 3: Solve the simplified program.
We note that within any interval $\alpha_{i,\ell},\alpha_{i,\ell+1})$, the order of $\eta$’s is preserved, that is, there exists a bijection $\pi_{i,\ell}:[n-1]\rightarrow A\setminus\{i\}$ such that $\eta_{\pi_{i,\ell}(1)}(\alpha,p(i))\leq\ldots\leq\eta_{\pi_{i,\ell}(n-1)}(\alpha,p(i))$. For convenience, we set $\eta_{\pi_{i,\ell}(0)}(\alpha,p(i))=$-$\infty$ and $\eta_{\pi_{i,\ell}(n)}(\alpha,p(i))=+\infty$.
$\min_{\alpha,p(i)}\$$\alpha f(i)+p(i)\cdot v(\{i\})+\eta_{\pi_{i,\ell}(k+1)}(\alpha,p(i))\cdot v(A\setminus\{i,\pi_{i,\ell}(1),\ldots,\pi_{i,\ell}(k)\})$
Let us observe that Program (3) is a constrained minimization problem in two variables, namely, $\alpha,p(i)$ which we are able to solve due to its specific structure. In particular, we utilize the fact that there are only two variables, and one of them only appears linearly in the objective function. More specifically, one can define $\beta=1/\alpha$. Then, the program is minimizing a function of the form $C_{1}/\beta+C_{2}\cdot p(i)+C_{3}\cdot\beta+C_{4}$ (where $C_{1},C_{2},C_{3},C_{4}$ are constants) subject to $\beta,p(i)$ belonging to a polygon with four sides (the four constraints). Since $p(i)$ only appears as a linear term in the objective function, we know that the optimum will be on the edges of the polygon. Thus, for each such edge, we rewrite the objective function as minimizing a function of one variable in the form of a quadratic numerator over a linear denominator. For this minimization, there is a solution in closed form.
Step 4: Output.
The algorithm considers all the inspection schemes in Steps 1, 2, 3 above for each $i\in A$, $\ell\in{n-1\choose 2}$, $k\in\{0\}\cup[n-1]$. For each, it computes, using Program (3), the optimal $\alpha,p(i)$, and calculates the optimal inspection scheme using Lemma 4.3 with $p(j)=0$ for $j\in\{\pi_{i,\ell}(1),\ldots,\pi_{i,\ell}(k)\},$ $p(j)=\eta_{j}(\alpha,p(i))$ for $j\in\{\pi_{i,\ell}(k+1),\ldots,\pi_{i,\ell}(n-1)\},$ and $p(i)$. It returns the inspection scheme with the maximal expected utility among all suggestions.